A coin is tossed from a height of 1.2 m from the ground. Assume that the coin follows a perfect vertical path with an initial velocity of 3.0 m/s.


a.How long should the coin travel to go back to its original position?

b.What is the velocity of the coin after 5.0 s?

c.What is its displacement after 0.52 s?


1
Expert's answer
2021-10-30T11:15:48-0400

H=(V2-U2)

1.2= 32-U2

1.2= 9-U2

1.2-9=-U2

-7.8=- U2

U= 2.79m